Dewey Corp. is expected to have an EBIT of $2,950,000 next year. Depreciation, the increase in net working capital, and capital spending are expected to be $230,000, $135,000, and $235,000, respectively. All are expected to grow at 18 percent per year for four years. The company currently has $18,000,000 in debt and 845,000 shares outstanding. After Year 5, the adjusted cash flow from assets is expected to grow at 3.1 percent indefinitely. The company’s WACC is 9.4 percent and the tax rate is 25 percent. What is the price per share of the company's stock?
Answer #1

Expected FCF, FCF1 = EBIT * (1 - Tax Rate) + Depreciation - Increase in NWC – Capital Spending
Expected FCF, FCF1 = $2,950,000 * (1 - 0.25) + $230,000 - $135,000 - $235,000
Expected FCF, FCF1 = $2,072,500

Growth rate for next 4 years is 18% and a constant growth rate (g) of 3.10% thereafter

FCF2 = $2,072,500 * 1.180 = $2,445,550
FCF3 = $2,445,550 * 1.180 = $2,885,749
FCF4 = $2,885,749 * 1.180 = $3,405,184
FCF5 = $3,405,184 * 1.180 = $4,018,117
FCF6 = $4,018,117 * 1.031 = $4,142,679

WACC = 9.40%

Horizon Value = FCF6 / (WACC - g)
Horizon Value = $4,142,679 / (0.0940 - 0.0310)
Horizon Value = $65,756,809.5238

Value of Firm = $2,072,500/1.094 + $2,445,550/1.094^2 + $2,885,749/1.094^3 + $3,405,184/1.094^4 + $4,018,117/1.094^5 + $65,756,809.5238/1.094^5
Value of Firm = $53,044,882.5186

Value of Equity = Value of Firm - Value of Debt
Value of Equity = $53,044,882.5186 - $18,000,000
Value of Equity = $35,044,882.5186

Stock Price = Value of Equity / Number of Shares Outstanding
Stock Price = $35,044,882.5186 / 845,000
Stock Price = $41.47